Manuel traveled 5 hours nonstop to mexico, a total of 293 miles. He took a train part of the way, which averaged 70mph, and then took a bus the remaining distance, which averaged 40mph. how long was manuel on the train?
Answer by mananth(16946)

Train x miles --------- 70 mph
Bus 293-x miles---------40 mph
time train + time bus = 5
t=d/r
x/70 + (293-x)/40 = 5
Multiply by LCD 2800
40x+70(293-x)= 5 * 2800
-30x+20510=14000
-30x=-6510
x=-6510/-30
x=217----------- train distance
balance bus distance
m.ananth@hotmail.ca